- to LTU green link, circuit repaired and both links is currently accepting traffic again.
- July 16th: Our fiber supplier switch fiber pair due to a degrade signal
- July 10th: An OTDR measurement was performed on the fiber stretch between Hanebol and Karlstad.
- July 8th: An amplifier was replaced in the transmission node in Karlstad.
- July 7th: An amplifier was replaced in the transmission node in Hanebol.
- July 4th: A transponder was replaced in Gävle due to intermittent errors.
- July 2nd: We had a slightly degraded signal in a DWDM in Karlstad on the west red network. Intrusive troubleshooting got performed on site. Fiber got cleaned and one suspicious 1db damper (actually 3db damping) on the rx got removed.
- July 2nd: A transponder was replaced du to intermittent errors
1 pcs 2d 03:09
1 pcs 02:39
2 pcs 2d 05:48
1 pcs 01:58
6 pcs 01:28
7 pcs 03:26